# Numerical simulation and analysis of mixing enhancement due to chaotic advection using an adaptive approach for approximating the dilution index

### v1

@@ Acknowledgments
Appendix A Analytical solution of the advective process in the pulsed source-sink system
The analytical solution of the flow defined in (1) is derived in [43, 41]. It is shown that due to the operation of a point sink/source a particle does move on a straight line towards/away from the sink/source. Thereby its squared distance to the sink/source over time $t$ is reduced/increased by $\lambda^{2}(t):=\frac{Q}{\pi}t$. Hence, while operating the sink for time $t$, all particles within the circle of radius $\lambda(t)$ are extracted from the domain. Vise versa, by operating the source for time $t$, the newly injected particles fill the circle with radius $\lambda(t)$ around the source as indicated in Fig. 1. For a source at location $\left(\begin{smallmatrix}0\\ 0\end{smallmatrix}\right)$ we can define the resulting flow by
@@ Appendix A Analytical solution of the advective process in the pulsed source-sink system
In the case of a sink at the same location $+\lambda^{2}$ is replaced by $-\lambda^{2}$. Assuming now arbitrary locations for sink and source, we can convert this configuration to a setup with sink and source at ±$\pm\left(\begin{smallmatrix}a\\ 0\end{smallmatrix}\right)$ by using a coordinate transformation. Multiplying the length scale by $\frac{1}{a}$ and using $\Lambda=\frac{\lambda}{a}$, we can further assume sink and source to be located at ±$\pm\left(\begin{smallmatrix}1\\ 0\end{smallmatrix}\right)$, without loss of generality [41]. Given a source at location $\left(\begin{smallmatrix}-1\\ 0\end{smallmatrix}\right)$ the flow in (21) is modified to
@@
For the reinjection procedure, we find that each particle $\left(\begin{smallmatrix}x\\ y\end{smallmatrix}\right)$ within the circle of radius $\Lambda$ around $\left(\begin{smallmatrix}1\\ 0\end{smallmatrix}\right)$, i.e., $\left(\begin{smallmatrix}x\\ y\end{smallmatrix}\right)\in B_{\Lambda}\left(\begin{smallmatrix}1\\ 0\end{smallmatrix}\right)$, is extracted from the domain after time $p\tau$ ($p\in[0,1]$). Following the first-in-first-out protocol the same particle is injected by the source at time $p\tau$ after the source started operating. The total time the particle moves with the flow remains $\tau$ and hence the distance to the source after its full stroke must be $\sqrt{\Lambda^{2}-\left(x^{2}+y^{2}\right)}$. We additionally assume, the injection angle to be $\pi-\theta$ with $\theta$ being the extraction angle. This leads to the map $M_{\text{inj}}:B_{\Lambda}\left(\begin{smallmatrix}1\\ 0\end{smallmatrix}\right)\rightarrow\mathbb{R}^{2}$
@@
Appendix B Analytical solution to the advective process in the reference system of the RPM flow
@@
The analytical solution of Hamiltonian system of the reference system is derived in [38]. Thereby source and sink are fixed at locations ±$\pm\left(\begin{smallmatrix}0\\ 1\end{smallmatrix}\right)$ while the rotation of the RPM flow is implemented by rotating the particles. As the flow is symmetrical to the $y$-axis, we only respect the positive side of the $x$-axis. The authors of [38] utilize the fact that the streamlines of the flow are given by the contour lines of the Hamiltonian $\psi$, plotted in Fig. 14. The position of each particle can be represented uniquely by its angle to the origin $\theta\in(-\frac{\pi}{2},\frac{\pi}{2})$ and its streamline $\psi\in(0,\frac{\pi}{2}]$. A representation in this coordinate system is convenient as for each particle $\psi$ is constant in time. They establish an expression for the advection time of a particle along a streamline until it reaches $\theta=0$. It is given by
$\begin{split}t_{\text{adv}}(\theta,\psi)=&\csc^{2}(\psi)\\ \times\Biggl{\{}&\cot(\psi)\arctan\left[\frac{\sin(\theta)\cot(\psi)}{\sqrt{1+\cos^{2}(\theta)\cot^{2}(\psi)}}\right]\\ &+\sin(\theta)\sqrt{1+\cos^{2}(\theta)\cot^{2}(\psi)}\\ &-|\cot(\psi)|(\theta+\cos(\theta)\sin(\theta))\Biggl{\}}.\end{split}$ (22)
@@
Both quantities are plotted in Fig. 15. Combining (22) and (23), the travel time of a particle until it reaches the sink is given by $t_{\text{adv}}(\theta,\psi)+\frac{T_{\text{res}}(\psi)}{2}$. If a particle does not get extracted within time $t$ it holds that
@@
$\begin{split}t_{\text{adv}}&(\theta(t),\psi_{0})+\frac{T_{\text{res}}(\psi_{0})}{2}=\\ &\left(t_{\text{adv}}(\theta_{0},\psi_{0})+\frac{T_{\text{res}}(\psi_{0})}{2}-t\right)\mod T_{\text{res}}(\psi_{0}).\end{split}$ (24)
Thereby, the modulo operator $\mod$ takes care of the extraction reinjection process. If the travel time of a particle $(\theta_{0},\psi_{0})$ to the sink is smaller than $t$, i.e., $t_{\text{adv}}(\theta_{0},\psi_{0})+\frac{T_{\text{res}}(\psi_{0})}{2}<t$, it gets extracted by the sink. The reinjection by the source is performed instantaneously onto the same streamline. The new travel time to the sink is given by $\left(t_{\text{adv}}(\theta_{0},\psi_{0})+\frac{T_{\text{res}}(\psi_{0})}{2}-t\right)\mod T_{\text{res}}(\psi_{0})$. We remark that the corresponding equation in [38] uses a plus sign in front of the $t$ on the right-hand side. As we can reproduce the experimental results in [38] using the corrected version of the equation given in (24), we can show that the sign change is only a typo in the formula. The numerical results of [38] are still valid. Solving (24) yields the flow $\Psi_{1}^{t}:\{\left(\begin{smallmatrix}x\\ y\end{smallmatrix}\right)\in B_{1}\left(\begin{smallmatrix}0\\ 0\end{smallmatrix}\right):x>0\}\rightarrow\{\left(\begin{smallmatrix}x\\ y\end{smallmatrix}\right)\in B_{1}\left(\begin{smallmatrix}0\\ 0\end{smallmatrix}\right):x>0\}$ which can be used for all particles on the positive side of the $x$-axis. For particles on the negative side, we use the reflection symmetry of $\psi$ with respect to the $y$-axis, yielding the flow $\Psi_{2}^{t}:\{\left(\begin{smallmatrix}x\\ y\end{smallmatrix}\right)\in B_{1}\left(\begin{smallmatrix}0\\ 0\end{smallmatrix}\right):x<0\}\rightarrow\{\left(\begin{smallmatrix}x\\ y\end{smallmatrix}\right)\in B_{1}\left(\begin{smallmatrix}0\\ 0\end{smallmatrix}\right):x<0\}$, $\Psi_{2}^{t}\left(\begin{smallmatrix}x\\ y\end{smallmatrix}\right)=\left(\begin{smallmatrix}-1&0\\ 0&1\end{smallmatrix}\right)\Psi_{1}^{t}\left(\begin{smallmatrix}-x\\ y\end{smallmatrix}\right)$.

### v2

@@ Acknowledgments
A Analytical solution of the advective process in the pulsed source-sink system
The analytical solution of the flow defined in (1) is derived in [45, 17]. It is shown that due to the operation of a point sink/source a particle does move on a straight line towards/away from the sink/source. Thereby its squared distance to the sink/source over time $t$ is reduced/increased by $\lambda^{2}(t):=Qt/pi$. Hence, while operating the sink for time $t$, all particles within the circle of radius $\lambda(t)$ are extracted from the domain. Vise versa, by operating the source for time $t$, the newly injected particles fill the circle with radius $\lambda(t)$ around the source as indicated in Fig. 1. For a source at location $\left(\begin{smallmatrix}0\\ 0\end{smallmatrix}\right)$ we can define the resulting flow by
@@ A Analytical solution of the advective process in the pulsed source-sink system
In the case of a sink at the same location $\lambda^{2}$ is replaced by $\lambda^{2}$. Assuming now arbitrary locations for sink and source, we can convert this configuration to a setup with sink and source at $\pm\left(\begin{smallmatrix}a\\ 0\end{smallmatrix}\right)$ by using a coordinate transformation. Multiplying the length scale by $1/a$ and using $\Lambda=\lambda/a$, we can further assume sink and source to be located at $\pm\left(\begin{smallmatrix}1\\ 0\end{smallmatrix}\right)$, without loss of generality [17]. Given a source at location $\left(\begin{smallmatrix}-1\\ 0\end{smallmatrix}\right)$ the flow in (21) is modified to
@@
For the reinjection procedure, we find that each particle $\left(\begin{smallmatrix}x\\ y\end{smallmatrix}\right)$ within the circle of radius $\Lambda$ around $\left(\begin{smallmatrix}1\\ 0\end{smallmatrix}\right)$, i.e., $\left(\begin{smallmatrix}x\\ y\end{smallmatrix}\right)\in B_{\Lambda}\left(\begin{smallmatrix}1\\ 0\end{smallmatrix}\right)$, is extracted from the domain after time $p\tau$ ($p\in[0,1]$). Following the first-out-first-in protocol the same particle is injected by the source at time $p\tau$ after the source started operating. The total time the particle moves with the flow remains $\tau$ and hence the distance to the source after its full stroke must be $\sqrt{\Lambda^{2}-\left(x^{2}+y^{2}\right)}$. We additionally assume, the injection angle to be $\pi-\theta$ with $\theta$ being the extraction angle. This leads to the map $M_{\text{inj}}:B_{\Lambda}\left(\begin{smallmatrix}1\\ 0\end{smallmatrix}\right)\rightarrow\mathbb{R}^{2}$
@@
B Analytical solution to the advective process in the reference system of the RPM flow
@@
The analytical solution of Hamiltonian system of the reference system is derived in [26]. Thereby source and sink are fixed at locations $\pm\left(\begin{smallmatrix}0\\ 1\end{smallmatrix}\right)$ while the rotation of the RPM flow is implemented by rotating the particles. As the flow is symmetrical to the $y$-axis, we only respect the positive side of the $x$-axis. The authors of [26] utilize the fact that the streamlines of the flow are given by the contour lines of the Hamiltonian $\psi$, plotted in Fig. 14. The position of each particle can be represented uniquely by its angle to the origin $\theta\in(-\pi/2,\pi/2)$ and its streamline $\psi\in(0,\pi/2]$. A representation in this coordinate system is convenient as for each particle $\psi$ is constant in time. They establish an expression for the advection time of a particle along a streamline until it reaches $\theta=0$. It is given by
$\begin{split}t_{\text{adv}}(\theta,\psi)=\csc^{2}(\psi)\times\Biggl\{&\cot(\psi)\arctan\left[\frac{\sin(\theta)\cot(\psi)}{\sqrt{1+\cos^{2}(\theta)\cot^{2}(\psi)}}\right]\\ &+\sin(\theta)\sqrt{1+\cos^{2}(\theta)\cot^{2}(\psi)}-|\cot(\psi)|(\theta+\cos(\theta)\sin(\theta))\Biggl\}.\end{split}$ (22)
@@
Both quantities are plotted in Fig. 15. Combining (22) and (23), the travel time of a particle until it reaches the sink is given by $t_{\text{adv}}(\theta,\psi)+T_{\text{res}}(\psi)/2$. If a particle does not get extracted within time $t$ it holds that
@@
$\begin{split}t_{\text{adv}}(\theta(t),\psi_{0})+\frac{T_{\text{res}}(\psi_{0})}{2}=\left(t_{\text{adv}}(\theta_{0},\psi_{0})+\frac{T_{\text{res}}(\psi_{0})}{2}-t\right)\mod T_{\text{res}}(\psi_{0}).\end{split}$ (24)
Thereby, the modulo operator $\mod$ takes care of the extraction reinjection process. If the travel time of a particle $(\theta_{0},\psi_{0})$ to the sink is smaller than $t$, i.e., $t_{\text{adv}}(\theta_{0},\psi_{0})+T_{\text{res}}(\psi_{0})/2<t$, it gets extracted by the sink. The reinjection by the source is performed instantaneously onto the same streamline. The new travel time to the sink is given by $\left(t_{\text{adv}}(\theta_{0},\psi_{0})+T_{\text{res}}(\psi_{0})/2-t\right)\mod T_{\text{res}}(\psi_{0})$. We remark that the corresponding equation in [26] uses a plus sign in front of the $t$ on the right-hand side. As we can reproduce the experimental results in [26] using the corrected version of the equation given in (24), we can show that the sign change is only a typo in the formula. The numerical results of [26] are still valid. Solving (24) yields the flow $\Psi_{1}^{t}:\{\left(\begin{smallmatrix}x\\ y\end{smallmatrix}\right)\in B_{1}\left(\begin{smallmatrix}0\\ 0\end{smallmatrix}\right):x>0\}\rightarrow\{\left(\begin{smallmatrix}x\\ y\end{smallmatrix}\right)\in B_{1}\left(\begin{smallmatrix}0\\ 0\end{smallmatrix}\right):x>0\}$ which can be used for all particles on the positive side of the $x$-axis. For particles on the negative side, we use the reflection symmetry of $\psi$ with respect to the $y$-axis, yielding the flow $\Psi_{2}^{t}:\{\left(\begin{smallmatrix}x\\ y\end{smallmatrix}\right)\in B_{1}\left(\begin{smallmatrix}0\\ 0\end{smallmatrix}\right):x<0\}\rightarrow\{\left(\begin{smallmatrix}x\\ y\end{smallmatrix}\right)\in B_{1}\left(\begin{smallmatrix}0\\ 0\end{smallmatrix}\right):x<0\}$, $\Psi_{2}^{t}\left(\begin{smallmatrix}x\\ y\end{smallmatrix}\right)=\left(\begin{smallmatrix}-1&0\\ 0&1\end{smallmatrix}\right)\Psi_{1}^{t}\left(\begin{smallmatrix}-x\\ y\end{smallmatrix}\right)$.
@@
Supplementary material
S1 Introduction
The supporting material shows our numerical results in more details. For the motivation of our adaptive approach of selecting the dilution index we show numerical results also for different amounts of particles, highlighting how the approximation error can be decreased by simultaneously increasing the amount of particles and decreasing the grid size in section Appendix S2. For the numerical experiment concerning mixing enhancement of the RPM flow we show the results for the dilution index, the growth rate and the convergence rate, for all considered parameter combinations, in section Appendix S3.
S2 A computationally reliable approach for the approximation of the dilution index
The usage of an increased number of particles in the particle tracking experiments enables the method to select smaller grid sizes for the dilution index approximation. Similar to Fig. 6 in the paper we created Fig. S1. Plot (a) shows the dilution index for different values of the diffusion coefficient $\sigma$ and different amounts of particles $m$. We show the upper limit $E\leq m\Delta V$ for the approximated dilution index for all considered choices of $m$. For a lager $m$ we find the plateau in Fig. S1(a) to be elongated compared to smaller $m$. We also show the relative approximation error in Fig. S1(b) and $\Delta_{\delta}[E](h)$ in Fig. S1(c). We find that, as expected, the approximation error of the dilution index $E$ reaches lower values the more particles are used in the particle tracking experiment. The observations made in the paper for m=10 000m=10\,000 are also valid for different choices of $m$.
S3 Numerical results for the RPM flow
In the paper we concentrate on the evolution of the reactor ratio $M$ over time $t$ for flow parametrizations $\left(\Theta,\tau\right)\in\left\{\left(0,0.2\right),\left(\pi/6,0.2\right),\left(\pi/6,0.5\right),\left(\pi/2,0.2\right),\left(\pi/2,0.5\right)\right\}$. The graph including all considered flow parametrizations is given in Fig. S2 and Fig. S3. We observe very different results depending on the different choices of the flow parametrization, due to the very individual KAM island structure.
Figure S1: (a) Shows the value of the dilution index $E$ over the grid size $h$ for different diffusion coefficients $\sigma$ and numbers of particles $m$. The dotted line gives the upper bound of the dilution index $m\Delta V$. The continuous lightened lines help to visualize the plateaus forming at the value of the analytical dilution index. (b) Shows the relative error $\frac{|\mathcal{E}-E|}{\mathcal{E}}$ of the approximated dilution index $E$ depending on the grid size $h$ to the analytical value $\mathcal{E}$. We observe a convergence $E\rightarrow\mathcal{E}$, until we pass the optimal grid size $h^{*}$. For smaller $h$, the dilution index dissipates from the analytical value. (c) Shows the approximated derivative of the dilution index $\Delta_{\delta}[E](h)$, using forward finite differences in the log-log scale, for $\sigma\neq 0$. The minima of the derivatives correspond approximately to the minima in the approximation error of plot (b).
Figure S2: Convergence of the reactor ration $M$ to $M_{\text{max}}=1$. (a)-(b) show the evolution of $M$ over $t$ for a logarithmic typed $y$-axis highlighting the convergence $M\rightarrow 1$, (c)-(d) show the convergence rate for $M$ approaching 11 for different flow parametrization. The convergence rate was computed using the moving average over the first order finite difference by using a window of $\pm 0.2$ in the exponent. The dotted lines correspond to the case without advection where the initial particle position is either at one point (gray) or in a ring (black).
Figure S3: (a)-(b) reactor ratio $M$ and (c)-(d) reactor ratio growth rate over time $t$ for different flow parametrization. The growth rate was approximated using the moving average over the first order finite differences with a window of $\pm 0.2$ in the exponent. The dotted lines correspond to the case without advection where the initial particle position is either at one point (gray) or in a ring (black).